OpenStudy (anonymous):

f(x) = xe^3/x a) Use l'Hospital's Rule to determine the limit as x 0+. b)Use calculus to find the local minimum value. c)Find the interval where the function is concave up. d) Find the interval where the function is concave down.

OpenStudy (anonymous):

e^3???

OpenStudy (anonymous):

r u sure thats right?

OpenStudy (anonymous):

x*e^(3/x)

OpenStudy (anonymous):

first limit is infinity

OpenStudy (anonymous):

ok i got that...

OpenStudy (anonymous):

assuming it is \[\lim_{x\rightarrow 0^+}xe^{\frac{3}{x}}\]

OpenStudy (anonymous):

derivative via product and chain rule gives \[\frac{e^{\frac{3}{x}} (x-3)}{x}\] yes?

OpenStudy (anonymous):

critical point at x = 3 (also 0 but not in the domain so can ignore that one

OpenStudy (anonymous):

so local min at x = 3. point is \[(3,3e)\]

OpenStudy (anonymous):

3 doing work

OpenStudy (anonymous):

neither does 3e

OpenStudy (anonymous):

i will guess concave up on \[(0,\infty)\][

OpenStudy (anonymous):

?

OpenStudy (anonymous):

critical point is x = 3, yes? not sure what you meant. did you mean "3 doesn't work?"

OpenStudy (anonymous):

i plugged it into my online homework...its wrong....I got the concave up and down already

OpenStudy (anonymous):

what does the homework say exactly? does it want both values? i mean both input and output? because i am fairly certain this is right

OpenStudy (anonymous):

just the local minimum

OpenStudy (anonymous):

well if that is what is says, the local minimum is \[3e\]

OpenStudy (anonymous):

\[\color{#ff0099}{\text{hello my myininaya!}}\]

OpenStudy (anonymous):

did i screw up somewhere? looks good to me

myininaya (myininaya):

\[f(x)=xe^{\frac{3}{x}}=>f'(x)=1*e^\frac{3}{x}+x*\frac{-3}{x^2}*e^\frac{3}{x}\] \[e^\frac{3}{x}(1+\frac{-3}{x})=0\] \[e^\frac{3}{x}\neq0, \frac{x-3}{x}=0=>x=3=>f(3)=3*e^\frac{3}{3}=3*e^1=3e\] you are right

OpenStudy (anonymous):

its 8.15

OpenStudy (anonymous):

i had to find the value for e

myininaya (myininaya):

3e=8.154845 approximately your question doesn't say round to the hundredths

OpenStudy (anonymous):

yes it does on the online problem thanks

myininaya (myininaya):

ok but you got it wrong thats weird

myininaya (myininaya):

or are you didn't round?

myininaya (myininaya):

i'm confused whats wrong then?

OpenStudy (anonymous):

i put 3e in